# Extension of Lipschitz maps definable in Hensel minimal structures

### v4

@@ Abstract.
In this paper, we establish a theorem on extension of Lipschitz maps $f$ definable in Hensel minimal, non-trivially valued fields $K$ of equicharacteristic zero. This may be regarded as a definable, non-Archimedean, non-locally compact version of Kirszbraun’s extension theorem. We proceed with double induction with respect to the dimensions of the ambient space and of the domain of $f$. To this end we introduce the concept of a definable open cell package with a skeleton, which along with the concept of a risometry plays a key role in our induction procedure.
@@ 3. Lipschitz extension on the affine line
The concept of a 0-definable reparametrized open cell is especially simply on the affine line (and only for dimension $1$). Indeed, the set of centers $c_{\xi}\in K$ of $C$ is of course finite, and consists of some $d$ distinct points $\{c_{1},\ldots,c_{d}\}\subset K$. Put
@@
Furthermore, the auxiliary sort $RV$ is stably embedded (cf. [4, Proposition 2.6.12]). We have even more:
Observation 3.2.
$\chi:K\to RV(K)^{t}$
with some $t\in\mathbb{N}$ such that, for any $a\in K$, the $\{a\}$-definable fiber $X_{a}\subset RV(K)^{n}$ is $\chi(a)$-definable.
This follows directly from the proof of loc.cit., and will be used in the proof of the extension theorem in the higher dimensional case as well. Hence each subset $C_{i}$ is a $\xi_{i}$-definable open cell for some $\xi_{i}\in RV(K)^{t}$.
We are now going to introduce the concept of a skeleton of $C$. Though this can be done in arbitrary Hensel minimal structures, with no additional conditions, we first suppose for simplicity that the infima
(3.2) $\rho(c_{i}):=\inf\,\{|x-c_{i}|:\ x\in C_{i}\}\in|K|$
are well defined for $i=1,\ldots,d$. In the general case where such infima do not exist, the reasoning below can be repeated verbatim with $\rho(c_{i})$, $i=1,\ldots,d$, replaced by the Dedekind cuts determined by the closed downward subsets of all elements of $|K|$ which are smaller than the sets
@@
Indeed, they only serve to group the centers of cells under study.
Remark 3.3.
For any Dedekind cut $\rho$ on the value group $|K|$ and an $r\in|K|$, the strong inequalities $r<\rho$ and $\rho<r$ have a clear meaning. The equality $\rho=r$ means that the cut $\rho$ represents the value $r$.
@@
take $r_{0}=0$ if 0 occurs among the $\rho(c_{i})$. For $i=0,\ldots,s$, $j=1,\ldots,t_{i}$, let $c_{i,j}$ be the centers with $\rho(c_{i,j})=r_{i}$; obviously, $s\leq d$ and
@@
We can canonically define, by modifying the initial centers, a skeleton $S(C)$ of $C$, which is a 0-definable set of centers satisfying the following metric condition:
(3.3) $\rho(c_{i,j})=r_{i}\ \ \text{and}\ \ |c_{i,j}-c_{p,k}|>r_{p}\ \ \text{for}\ \ i\leq p.$
It will be built through successive steps taking into account the values $r_{i}$ considered above.
@@
for some $k$. Note that, for any such removed center $c_{1,j}$, the cell $C_{1,j}$ is also a cell with center $c_{0,k}$ (with the same set $R_{1,j}$).
Next notice that the equivalence relation
@@
lies in no cell $C_{1,j_{1}},\ldots,C_{1,j_{l}}$, or it lies in a ball of radius $r_{1}$ from exactly one of those cells, say $C_{1,j_{k}}$. We can thus replace each center $c_{1,j_{q}}$ from such a maximal set with $c$ or $c_{1,j_{k}}$ (without change of the set $R_{1,j_{q}}$), according as the former or latter condition holds, and add such a new center to the skeleton. In this manner, any such equivalence class determines one point to be added to the skeleton. By Remark 3.4, the set of centers from the skeleton, obtained at this level 1, form a 0-definable set satisfying condition 3.3 for $0\leq i\leq 1$.
@@
$\bullet$ Further, repeat the above construction at level 3 for the centers $c_{3,j}$, and so on. This canonical process leads eventually to a unique 0-definable set of centers $S(C)=\{c_{1},\ldots,c_{e}\}$ with $e\leq d$, satisfying condition 3.3, sometimes regarded as a $d$-tuple too. Then we say that the reparametrized cell $C$ is with a skeleton $(C)$. It is easy to see that $S(C)\cap C=\emptyset$ and that center exchange in the construction of skeleton does not change the distance from the center.
@@
$(c_{1},R_{1}),\ldots,(c_{e},R_{e}).$
Remark 3.5.
In view of the above, the extension problem on the affine line comes down to extending Lipschitz 0-definable functions from such a finitary 0-definable open cell package.
Remark 3.6.
Observe that the skeleton $S(C)$ of a reparametrized cell $C$ depends on the initial set of centers of $C$. Furthermore, there are finitely many configurations, say
$\mathfrak{C}_{1},\ldots,\mathfrak{C}_{N},\ \ N=N(d),$
of the initial $d$-tuples $(c_{1},\ldots,c_{d})$ of centers $c_{i}$ or, more precisely, of $d$-tuples of pairs $(c_{i},\rho(c_{i}))_{i=1}^{d}$, such that for each of which the skeleton is given by a unique formula (built from the $d$-tuples and some arithmetic averages after a suitable choice, as described above) depending only on a given configuration.
In the sequel, we shall still need the following lemma whose proof uses the concept of a skeleton.
@@
Let $\mathcal{B}$ be a 0-definable family of disjoint open balls in $K$ which satisfies the following condition: for all balls $B_{1},B_{2}\in\mathcal{B}$, $B_{1}\neq B_{2}$, with radii $r_{1},r_{2}\in|K|\setminus\{0\}$, respectively, we have
@@ 4. Lipschitz extension on the affine plane
and next the function $\tilde{f}_{B_{1}}$ by gluing the functions $\tilde{f}_{B_{1},1},\ldots,\tilde{f}_{B_{1},p}$ just constructed. Eventually, we can canonically define a function $\tilde{f}_{1}$ by gluing the functions $\tilde{f}_{B_{1}}$ where $B_{1}$ runs over all maximal sets that satisfy condition $\Lambda_{1}$.
It is not difficult to check that all the functions $\tilde{f}_{B_{1}}$ are 1-Lipschitz on the $\lambda_{1}$-neighbourhood of the finite set
@@
and of a (finite) set of dimension <1<1, we can assume that $D=\mathrm{int}(D_{d})$, and thus, as before in Case II, that the constant cardinality of the fibers of $p$ is $d$. Therefore the 0-definable family
@@
$D_{\lambda}=\bigcup_{\lambda\in\bar{\xi}}\,D_{\bar{\xi}}$
for $\lambda\in\sigma(C)$. At the cost of a further finite 0-definable partitioning of $D$, we can also require that all (ordered) fibers
$(c_{\xi_{1},2}(x_{1}),\ldots,c_{\xi_{d},2}(x_{1}))$
of $p$, regarded as $d$-tuples, have the same configuration with respect to construction of a skeleton on the affine line.
Next, we wish to find a certain special 0-definable cell decomposition of $D$ which is finer than the covering $\{D_{\bar{\xi}}:\ {\bar{\xi}}\in\sigma(C)^{d}\}$. To this end consider a 0-definable family
$\Lambda_{a}:=\{\bar{\xi}\in\sigma(C)^{d}:\,a\in D_{\bar{\xi}}\},\ \ \ a\in D,$
of subsets of $\sigma(C)^{d}$. What will be used here is the fact that the auxiliary sort $RV$ is stably embedded (cf. [4, Proposition 2.6.12]). By Observation 3.2, there exists a 0-definable map
$\chi:K\to RV(K)^{s}$
such that each set $\Lambda_{a}$ is $\chi(a)$-definable and the family
$\Lambda_{\chi(a)}:=\Lambda_{a},\ \ \ \chi(a)\in\chi(D),$
is well defined and 0-definable, too, because $\Lambda_{a}$ is defined by a formula $\varphi(\bar{\xi},\chi(a))$. Therefore the 0-definable equivalence relation
$R(\chi(a),\chi(b))\ \ \Longleftrightarrow\ \ \Lambda_{\chi(a)}=\Lambda_{\chi(b)},\ \ a,b\in D,$
allows us to code the same fibres from among $\Lambda_{\chi(a)}$ with $\chi(a)\in\chi(D)$, by the imaginary elements (equivalence classes) induced by $R$:
$\ulcorner\Lambda_{\chi(a)}\urcorner=[\chi(a)]_{R}=[\chi(a)].$
In other words, we can parametrize the same fibres from among $\Lambda_{\chi(a)}$ by the imaginary elements $\chi(a)]$ with $\chi(a)\in\chi(D)$. Consider the two 0-definable families
$S_{a}=S_{\chi(a)}=S_{[\chi(a)]}:=\bigcap_{a\in D_{\bar{\xi}}}\,D_{\bar{\xi}}=\bigcap_{\bar{\xi}\in\Lambda_{a}}\,D_{\bar{\xi}}=\bigcap_{\bar{\xi}\in\ulcorner\Lambda_{\chi(a)}\urcorner}\,D_{\bar{\xi}}$
@@
$T_{a}=T_{\chi(a)}=T_{[\chi(a)]}:=\left\{b\in S_{[\chi(a)]}:\ \Lambda_{\chi(b)}=\Lambda_{\chi(a)}\right\}$
parametrized by $\chi(a)\in\chi(D)$, and also by the imaginary elements $\chi(a)]\in\chi(D)/R$. The second family is of course a partition of $D$.
Now, punctual construction of a skeleton from Section 3, performed here over parameters $x_{1}\in D$, leads to $e$ new centers
$c_{1,2}(x_{1}),\ldots,c_{e,2}(x_{1})$
with a number $e\leq d$ independent of $x_{1}$, which are produced from the initial $d$ centers $c_{\xi_{1}},\ldots,c_{\xi_{d}}$ by the same formulae. We call the graphs of those new centers (or their union) a skeleton of $C$ at level 2 (i.e. with respect to the variable $x_{2}$). Those centers are 1-Lipschitz on each set $T_{a}=T_{\chi(a)}=T_{[\chi(a)]}$. Indeed, each $c_{j,2}$ over each set $T_{\chi(a)}$ is built from the centers $\bar{\xi}$ for every $\bar{\xi}\in\Lambda_{a}$ as one of them or an average (construction of a skeleton on the affine line). Furthermore, the part $C\cap\pi_{1}^{-1}(T_{[\chi(a)]}$ of $C$ over each $T_{[\chi(a)]}$ is the union of $e$ open sets
$\bigcup_{j=1}^{e}\,\{x\in T_{[\chi(a)]}\times K:\ (rv(x_{1}-c_{\xi,1})),rv(x_{2}-c_{j,2}(x_{1}))\in R_{[\chi(a)],j}\},$
where $R_{[\chi(a)],j}$ is the union of all $R_{\xi}$ with $c_{\xi,2}(a)=c_{j,2}(a)$. We call it an open cell package determined by those centers and that family.
It follows from Theorem 2.1 that $D$ is a finite disjoint union of 0-definable reparametrized Lipschitz cells $D_{l}$ compatible with the family $T_{\chi(a)}$. By the LE-property and the induction hypothesis on the dimension $k$ of the subset $A$, we can assume, throwing away pieces of lower dimension, that
$D=\bigcup\,\left\{D_{\eta}:\ \eta\in\tau(D)\right\}\subset K_{x_{1}},$
is one of those open cells. Then each set $T_{a}=T_{[\chi(a)]}$ is a union of some twisted boxes (here balls) of $D$. We can also assume that $D$ is a 0-definable package determined by a family
@@
Summing up, for each $\eta$, the skeleton consists of the same number of graphs of those new $e$ 1-Lipschitz centers over $D_{\eta}$, which are not $\eta$-definable individually, but their union is $\eta$-definable.
Further, we have the following presentation
$C=\bigcup_{i=1}^{q}\bigcup_{j=1}^{e}\,C_{i,j},\ \ \ C_{i,j}:=\bigcup_{\lambda_{1}\in R_{i}}\,C_{i,j,\eta},$
@@
where
$R_{i,j,\lambda_{1}}:=\{\lambda_{2}\in G(K):\ \text{for some }\xi\ \text{one has:}$
$c_{\xi,1}=c_{i,1},\ c_{\xi,2}|D_{\lambda_{1}}=c_{j,2}|D_{\lambda_{1}},\ (\lambda_{1},\lambda_{2})\in R_{\xi}\}.$
Then we shall say that $C$ is a 0-definable open cell package determined by the finitary skeleton
@@
$(R_{i},R_{i,j,\lambda_{1}}),\ \ i=1,\ldots,q,\ j=1,\ldots,e,\ \lambda_{1}\in R_{i},$
expressed equivalently by the 0-definable family
$R_{i,j}\subset G(K)^{2},\ \ i=1,\ldots,q,\ j=1,\ldots,e,$
with
$R_{i,j}:=\{\lambda\in G(K)^{2}:\ \lambda_{2}\in R_{i,j,\lambda_{1}}\}.$
Then we get
@@
Furthermore, each center $c_{j,2}$ is 1-Lipschitz on every ball $D_{\eta}$.

### v5

@@ Abstract.
In this paper, we establish a theorem on extension of Lipschitz maps $f$ definable in Hensel minimal, non-trivially valued fields $K$ of equicharacteristic zero. This may be regarded as a definable, non-Archimedean, non-locally compact version of Kirszbraun’s extension theorem. We shall proceed with double induction with respect to the dimensions of the ambient space and of the domain of $f$. To this end, we introduce the concept of a definable open cell package with a skeleton which, along with the concept of a risometry, plays a key role in our induction procedure.
@@ 3. Lipschitz extension on the affine line
The concept of a 0-definable reparametrized open cell on the affine line is quite simple in comparison with higher dimensions. Indeed, the set of centers $c_{\xi}\in K$ of $C$ is finite, and consists of some $d$ distinct points $\{c_{1},\ldots,c_{d}\}\subset K$. Put
@@
We are now going to introduce the concept of a skeleton of $C$. The infima
(3.2) $\rho(c_{i}):=\inf\,\{|x-c_{i}|:\ x\in C_{i}\}\in|K|,\ \ i=1,\ldots,d,$
do not exist in general. Instead we shall consider the Dedekind cuts determined by the closed downward subsets of all elements of $|K|$ which are smaller than the sets
@@
They only serve to group the centers of the cells under study. For any Dedekind cut $\rho$ on the value group $|K|$ and an $r\in|K|$, the inequalities $r<\rho$ and $r>\rho$ have a clear meaning.
@@
take $r_{0}=0$ if 0 occurs among the $\rho(c_{i})$. For $i=0,\ldots,s$, let
$c_{i,j}\ \ \text{with}\ \ j=1,\ldots,t_{i},$
be the centers for which $\rho(c_{i,j})=r_{i}$; obviously, $s\leq d$ and
@@
We can canonically define, by modifying the initial centers, a skeleton $S(C)$ of $C$, which is a 0-definable set of centers satisfying the following elementary metric condition:
(3.3) $|c_{i,j}-c_{p,k}|>r_{i}\ \text{and}\ |c_{i,j}-c_{p,k}|>r_{p},$
which plays a crucial role in Lipschitz extension.
Skeleton will be built through successive steps taking into account the Dedekind cuts $r_{i}$ considered above. Now we outline its construction.
@@
for some $k$. Note that, for any such removed center $c_{1,j}$, the cell $C_{1,j}$ is also a cell with center $c_{0,k}$ (with the same set $R_{1,j}$). Next notice that the equivalence relation
@@
lies in no cell $C_{1,j_{1}},\ldots,C_{1,j_{l}}$, or it lies in a ball of radius $r_{1}$ from exactly one of those cells, say $C_{1,j_{k}}$. We can thus replace each center $c_{1,j_{q}}$ from such a maximal set with $c$ or $c_{1,j_{k}}$ (without change of the set $R_{1,j_{q}}$), according as the former or latter condition holds, and add such a new center to the skeleton. In this manner, any such equivalence class determines one point to be added to the skeleton. By Remark 3.2, the set of centers from the skeleton, obtained at this level 1, form a 0-definable set satisfying condition 3.3 for $0\leq i,p\leq 1$.
@@
$\bullet$ Further, repeat the above construction at level 3 for the centers $c_{3,j}$, and so on.
This canonical process leads eventually to a unique 0-definable set of centers
$S(C)=\{\tilde{c}_{1},\ldots,\tilde{c}_{e}\}\ \ \text{with}\ e\leq d,$
satisfying condition 3.3. Then we say that the reparametrized cell $C$ is with the skeleton $S(C)$. Obviously, the center exchange in this construction does not change the distance from the center to its balls.
@@
$(\tilde{c}_{1},R_{1}),\ldots,(\tilde{c}_{e},R_{e}).$
Observation 3.3.
Under the above assumptions, the skeleton $S(C)$ of a reparametrized open cell $C$ depends canonically on the initial set of centers of $C$. There are finitely many configurations, say $\mathfrak{C}_{1},\ldots,\mathfrak{C}_{N}$, $N=N(d)$, of the initial $d$-tuples $(c_{1},\ldots,c_{d})$ of centers, or rather, of $d$-tuples of pairs $(c_{i},\rho(c_{i}))_{i=1}^{d}$, such that for each of which the skeleton is given by a unique formula (built from the $d$-tuples and some arithmetic means occurring in the construction), which depends only on a given configuration.
It is not difficult to check that a single reparametrized open cell $C$, after such a canonical 0-definable modification of the parameter set $\sigma(C)$, can be treated as one whose set of centers is just its skeleton. This applies uniformly to definable families of reparametrized open cells with a common configuration.
The sort $RV$ is stably embedded (cf. [4, Proposition 2.6.12]). We have even more (which follows directly from the proof there):
Lemma 3.4.
$\chi:K\to RV(K)^{l},$
with some $l\in\mathbb{N}$, such that the $\{a\}$-definable fiber $X_{a}\subset RV(K)^{n}$ is $\chi(a)$-definable for any $a\in K$.
It should also be pointed out that, with the above notation, the pairs with Dedekind cuts
$(\chi(c_{i}),\rho(c_{i})),\ \ i=1,\ldots,d,$
are imaginary elements of the auxiliary sort $RV$. Further, dealing with a 0-definable family of cells with a valued field parameter, the foregoing configurations can be characterized in terms of certain ordinary and imaginary (Dedekind cuts) elements of $RV$, derived naturally from the initial cells of the family, by a 0-definable equivalence relation. Hence and by the above lemma, we obtain yet another
Observation 3.5.
Consider 0-definable families of open cells $C_{t}$ and of $d$-tuples $\bar{c}(t)$ of their centers on the affine line $K$, with a valued field parameter $t$. Then the induced family of configurations $\mathfrak{C}(C_{t},\bar{c}(t))$ can be coded by a certain imaginary parameter of the auxiliary sort $RV$. This will be used in the construction of a skeleton in the higher dimensional case; more precisely, to certain cell decompositions compatible with some $RV$-parameters.
By the LE-property (corollary 2.3), the extension problem on the affine line comes down to extending Lipschitz 0-definable functions from such a 0-definable open cell package. We shall still need a basic, though elementary, proposition whose proof uses the concept of a skeleton.
@@
Let $\mathcal{B}$ be a 0-definable family of pairwise disjoint open balls in $K$ which satisfies the following condition: for all balls $B_{1},B_{2}\in\mathcal{B}$, $B_{1}\neq B_{2}$, with radii $r_{1},r_{2}\in|K|\setminus\{0\}$, respectively, we have
@@ 4. Lipschitz extension on the affine plane
and next the function $\tilde{f}_{B_{1}}$ by gluing the functions $\tilde{f}_{B_{1},1},\ldots,\tilde{f}_{B_{1},p}$ just constructed. Eventually, we can canonically define a function $\tilde{f}_{1}$ by gluing the functions $\tilde{f}_{B_{1}}$ where $B_{1}$ runs over all maximal sets that satisfy condition $\Lambda_{1}$. It is not difficult to check that all the functions $\tilde{f}_{B_{1}}$ are 1-Lipschitz on the $\lambda_{1}$-neighbourhood of the finite set
@@
and of a (finite) set of dimension <1<1, we can assume that $D=\mathrm{int}(D_{d})$, and thus, as before in Case II, that the constant cardinality of the fibers $p^{-1}(x_{1})=C_{x_{1}}$ is $d$ for all $x_{1}\in D$. Then the 0-definable family
@@
$D_{\lambda}=\bigcup_{\lambda\in\bar{\xi}}\,D_{\bar{\xi}}\ \ \text{for every}\ \ \lambda\in\sigma(C).$
Next, we wish to find a certain special 0-definable cell decomposition of $D$, which is finer than the covering $\{D_{\bar{\xi}}:\ {\bar{\xi}}\in\sigma(C)^{d}\}$. To this end, consider the following two 0-definable families $\Lambda$ and $\Theta$ with parameter $a\in D$ and fibres in the auxiliary sort $RV$:
$\Lambda_{a}:=\{\bar{\xi}\in\sigma(C)^{d}:\,a\in D_{\bar{\xi}}\}\subset\sigma(C)^{d}$
@@
$\Theta_{a}:=\{(\bar{\xi},rv(a-c_{\xi_{1}}),\dots,rv(a-c_{\xi_{d}})):\,\bar{\xi}=(\xi_{1},\ldots,\xi_{d})\in\Lambda_{a}\}$
$\subset\sigma(C)^{d}\times G(K)^{d}.$
Next define the following 0-definable equivalence relation on $D$:
[…]
which induces the 0-definable partition of $D$ into its equivalence classes
$T_{a}=\{b\in D:\ \Theta_{a}=\Theta_{b}\},\ \ a\in D.$
It follows immediately from Theorem 2.1 and Observation 3.5 that there exists a finite 0-definable decomposition of $D$ into reparametrized Lipschitz cells, which is compatible both with the above partition and with the configurations $\mathfrak{C}(C_{a},\bar{\xi}(a))$, i.e. both the fibers $\Theta_{a}$ and all the configurations $\mathfrak{C}(C_{a},\bar{\xi}(a))$ with respect to the construction of a skeleton on the affine line and with $\bar{\xi}(a)\in\Lambda_{a}$, are constant over each twisted box $B$ (here, on the affine line, ball) of those cells.
By the LE-property and induction hypothesis, throwing away pieces of lower dimension, we can assume that $D$ is one of those reparametrized open cell. It follows immediately from the very construction that each such ball $B$ lies in a unique equivalence class $T_{a}$, and that the centers $c_{\xi_{i},1}(x_{1})$ are 1-Lipschitz on $B$ for every $\bar{\xi}=(\xi_{1},\ldots,\xi_{d})\in\Lambda_{a}$. Hence and by the constancy of fiber configuration, the punctual construction of a skeleton from Section 3, performed here over parameters $x_{1}\in B$, leads to $e$ new centers
$\tilde{c}_{1,2}(x_{1}),\ldots,\tilde{c}_{e,2}(x_{1})$
with a number $e\leq d$ independent of $x_{1}$, which are produced from the initial $d$ centers by the same formulae and which are 1-Lipschitz too. Indeed, the new centers of the skeleton are built as arithmetic means of a number of the initial ones.
We call the graphs of those new centers (or their union) a skeleton of $C$ at level 2 (i.e. with respect to the variable $x_{2}$). Similarly as before on the affine line, we can actually assume that the reparametrized cell $C$ has fibers just of constant cardinality $e$ over $D$. For simplicity, we drop the tilde sign.
Further, we can also assume that $D$ itself is a 0-definable package determined by a family
@@
Summing up, the skeleton consists of the same number $e$ of graphs of new centers that are 1-Lipschitz over each twisted box (here, on the affine line, ball) of the package $D$. We thus get the following presentation
$C=\bigcup_{i=1}^{q}\bigcup_{j=1}^{e}\,C_{i,j},\ \ \ C_{i,j}:=\bigcup_{\lambda_{1}\in R_{i}}\,C_{i,j,\lambda_{1}},$
@@
where $R_{i,j,\lambda_{1}}$ is defined by the formula
$\{\lambda_{2}\in G(K):\ \ \exists\;a\in D\ \exists\;\xi\in\bar{\xi}\in\Lambda_{a}\ \ [rv(a-c_{i,1})=\lambda_{1},\ (\lambda_{1},\lambda_{2})\in R_{\xi}]\};$
independence from the point $a$ follows directly from that the package $D$ is compatible with the partition $\Theta$.
We shall then say that the reparametrized cell $C$ is a 0-definable open cell package determined by the finitary skeleton
@@
$(R_{i},R_{i,j,\lambda_{1}}),\ \ i=1,\ldots,q,\ j=1,\ldots,e,\ \lambda_{1}\in R_{i};$
the latter can be expressed equivalently by the 0-definable family
$R_{i,j}:=\{\lambda=(\lambda_{1},\lambda_{2})\in G(K)^{2}:\ \lambda_{2}\in R_{i,j,\lambda_{1}}\}\subset G(K)^{2}$
with $i=1,\ldots,q$, $j=1,\ldots,e$. Hence we get
@@
Furthermore, each center $c_{j,2}$ is 1-Lipschitz on every twisted box (here ball) of the package $D$.